OpenStudy (anonymous):

Which two points are on the graph of G(e) = 40 – 8e? Choose exactly two answers that are correct. (Points : 1) (0, –40) (0, 40) (–5, 0) (5, 0) Question 2.2. Which two points are on the graph of ? Choose exactly two answers that are correct. (Points : 1) (0, –10) (0, 10) (4, 0) (25, 0)

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

but i need one more for it

OpenStudy (anonymous):

i meant the first one, i had to mark 2 boxes

OpenStudy (aum):

G(e) = 40 – 8e Test each point. (0, –40): put e = 0: G(e) = 40 - 0 = 40 \(\ne\) -40. False. (0, 40): put e = 0: G(e) = 40 - 0 = 40. True. (–5, 0): put e = -5: G(e) = 40 - 8 * (-5) = 40 + 40 = 80 \(\ne\) 0. False. (5, 0): put e = 5: G(e) = 40 - 8 * 5 = 40 - 40 = 0. True.

OpenStudy (anonymous):

can you help me with the second one? i posted the link above

OpenStudy (aum):

F(e) = 10 - 2/5 * e Try each point and see which one works. I will do the first answer choice. (0, –10) : put e = 0 and see if F(e) is -10: F(e) = 10 - 2/5 * 0 = 10 - 0 = 10 \(\ne\) -10. False. You try the others.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you tell me if my answer is correct?

OpenStudy (anonymous):

wait i changed my answer to 22.6

OpenStudy (aum):

(308 - 195) / (14 - 9) = 22.6 is correct.

OpenStudy (anonymous):

thanks i got a 100% on the test